Need help please.
6) The volume of a cube is given by V = s³. Find the length of a side of a cube if the Volume is 1000 cm³.
Answer:
Show work in this space.

Answer by fractalier(6550)

Here volume V = s^3, but we are given a volume of 1000 cubic cm. Thus
V = s^3 = 1000 and by taking the cube root of both sides, we get
s = 10 cm